Evaluate the difference quotient and simplify the result. 1 f(x) X + 5 f(x + Ax) – f(x) ΔΧ
Answer #1

f\left ( x \right )=\frac{1}{x+5}

\frac{f\left ( x+\Delta x \right )-f\left ( x \right )}{\Delta x}=\frac{\frac{1}{x+\Delta x+5}-\frac{1}{x+5}}{\Delta x}

=\frac{\frac{x+5-\left ( x+\Delta x+5 \right )}{\left (x+\Delta x+5 \right )\left ( x+5 \right )}}{\Delta x}

=\frac{\frac{x+5- x-\Delta x-5}{\left (x+\Delta x+5 \right )\left ( x+5 \right )}}{\Delta x}

=\frac{\frac{-\Delta x}{\left (x+\Delta x+5 \right )\left ( x+5 \right )}}{\Delta x}

=\frac{-\Delta x}{\left (x+\Delta x+5 \right )\left ( x+5 \right )\Delta x}

=\frac{-1}{\left (x+\Delta x+5 \right )\left ( x+5 \right )}

That is

\boldsymbol{\frac{f\left ( x+\Delta x \right )-f\left ( x \right )}{\Delta x}=\frac{-1}{\left (x+\Delta x+5 \right )\left ( x+5 \right )}}

-----------------------------------------------------------------